OpenStudy (fran345):

Can someone help me with these questions? Write the equation of the line, in standard form, that has a y-intercept of 2 and is parallel to 2x + y = -5. Write the equation of the line, in standard form, that has an x-intercept of 2 and is parallel to 2x + y = -5.

OpenStudy (3mar):

May I help?

OpenStudy (fran345):

yes

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

1 min

OpenStudy (3mar):

Write the equation of the line, in standard form, that has a y-intercept of 2 and is parallel to 2x + y = -5. Hint: All parallel lines have the same slope. So you got the slope and have the point, then you would use this:

OpenStudy (fran345):

the slope is -5?

OpenStudy (3mar):

For the 1st, the slope would be m=-2

OpenStudy (3mar):

2x + y = -5 y=-2x-5 m=-2 m is the coefficient of x.

OpenStudy (fran345):

I still don't know how you got -2. I know how to do it if I had the points with it's coordinates.

OpenStudy (3mar):

The slope-intercept form of a line is